# Groups of type $FP$ via graphical small cancellation

### 14 Dec 2023 v5

@@ 7 A spectacular 2-complex
Let $g$ be an element of $G$ of order $d$. Our complex will depend on the pair $(d,q)$ and on the conjugacy class of the element $g$. The centralizer in $G$ of $g$ is cyclic of order $q+\epsilon$, and the normalizer of the subgroup generated by $g$ is dihedral of order $2(q+\epsilon)$. It follows that the conjugacy class of $g$ contains $q(q-\epsilon)$ elements and that it is closed under taking inverses. The element $g$ is a power of an element of order $q+\epsilon$. From this it follows that the permutation action of $g$ on $K_{1}^{0}$ has $1-\epsilon$ fixed points and $(q+\epsilon)/d$ cycles of length $d$. We use each of these cycles of length $d$ to describe the attaching map for a $d$-gon. If $x$ is an element of the conjugacy class of $g$ and the vertices $v_{0},\ldots,v_{d-1},v_{d}$ are such that $v_{i}=x^{i}(v_{0})$, then there is a $d$-gon whose boundary is the edge loop of length $d$ in $K_{1}^{1}$ consisting of the edges $(v_{0},v_{1}),(v_{1},v_{2}),\ldots,(v_{d-1},v_{d})$. This edge loop passes through the vertices $v_{0},\ldots,v_{d-1}$ consecutively. Note that the vertex orbits for $x^{-1}$ give rise to the same polygons as the vertex orbits for $x$, with the opposite orientation. Each pair of the form $x,x^{-1}$ of elements of the conjugacy class of $g$ gives a recipe for attaching $(q+\epsilon)/d$ distinct $d$-sided polygons to $K_{1}^{1}$. The complex $K_{1}$ is thus defined by attaching $|G|/2d$ distinct $d$-gons to $K_{1}^{1}$.
@@ 8 Closing remarks
It is easy to see that the groups $H(S)$ fall into uncountably many quasi-isometry classes using Bowditch’s argument [4]. It is less clear what happens with the groups $G(S)$, except in the case when $G(S)$ isomorphic to a generalized Bestvina–Brady group $G_{L}(S)$, which was covered in [22]. Since this article was first submitted, the first named author has resolved the case when $Z=\{k^{n}\,\,:\,\,n\geq 0\}$ and $G_{P}$ is as described in Proposition 5.2, in [9]. The general case (i.e., for arbitrary $Z$ and embeddings $H_{P}\rightarrow G_{P}$) remains open. The methods introduced in [28] may prove useful but they do not seem to apply directly. In the language of [28], $\psi:S\mapsto G(S)$ defines an injective function from the set of subsets of $Z$ to the space of marked groups generated by the elements of $L$. To apply the main theorem of [28] it would suffice to show that the image of $\psi$ is closed, which in turn would follow if one could show that $\psi$ is continuous for the Tychonoff (or product) topology on the powerset of $Z$.

### 07 Feb 2024 v6

@@ 7 A spectacular 2-complex
Let $g$ be an element of $G$ of order $d$. Our complex will depend on the pair $(d,q)$ and on the conjugacy class of the element $g$. By Theorem 6.2, the centralizer in $G$ of $g$ is cyclic of order $q+\epsilon$, and the normalizer of the subgroup generated by $g$ is dihedral of order $2(q+\epsilon)$. It follows that the conjugacy class of $g$ contains $q(q-\epsilon)$ elements and that it is closed under taking inverses. The element $g$ is a power of an element of order $q+\epsilon$. From this it follows that the permutation action of $g$ on $K_{1}^{0}$ has $1-\epsilon$ fixed points and $(q+\epsilon)/d$ cycles of length $d$. We use each of these cycles of length $d$ to describe the attaching map for a $d$-gon. If $x$ is an element of the conjugacy class of $g$ and the vertices $v_{0},\ldots,v_{d-1},v_{d}$ are such that $v_{i}=x^{i}(v_{0})$, then there is a $d$-gon whose boundary is the edge loop of length $d$ in $K_{1}^{1}$ consisting of the edges $(v_{0},v_{1}),(v_{1},v_{2}),\ldots,(v_{d-1},v_{d})$. This edge loop passes through the vertices $v_{0},\ldots,v_{d-1}$ consecutively. Note that the vertex orbits for $x^{-1}$ give rise to the same polygons as the vertex orbits for $x$, with the opposite orientation. Each pair of the form $x,x^{-1}$ of elements of the conjugacy class of $g$ gives a recipe for attaching $(q+\epsilon)/d$ distinct $d$-sided polygons to $K_{1}^{1}$. The complex $K_{1}$ is thus defined by attaching $|G|/2d$ distinct $d$-gons to $K_{1}^{1}$.
@@ 8 Closing remarks
It is easy to see that the groups $H(S)$ fall into uncountably many quasi-isometry classes using Bowditch’s argument [4]. It is less clear what happens with the groups $G(S)$, except in the case when $G(S)$ isomorphic to a generalized Bestvina–Brady group $G_{L}(S)$, which was covered in [22]. Since this article was first submitted, the first named author has resolved the case when $Z=\{k^{n}\,\,:\,\,n\geq 0\}$ and $G_{P}$ is as described in Proposition 5.2, in [9]. The question of whether the groups $G(S)$ form uncountably many quasi-isomorphism classes for arbitrary recursively enumerable $Z\subseteq{\mathbb{Z}}$ and arbitrary choices of embeddings $H_{p}\rightarrow G_{P}$ remains open in full generality. The methods introduced in [28] may prove useful in resolving this but they do not seem to apply directly. Fix a finite generating set for $G(\emptyset)$, and use the image of this set under the surjective homomorphism $G(\emptyset)\rightarrow G(S)$ as a set of generators for $G(S)$. In the language of [28], the map $\psi:S\mapsto G(S)$ defines an injective function from the set of subsets of $Z$ to the space of marked groups. To apply the main theorem of [28] it would suffice to show that the image of $\psi$ is closed, which in turn would follow if one could show that $\psi$ is continuous for the Tychonoff (or product) topology on the powerset of $Z$.